# Trusted Unified Feature-Neighborhood Dynamics for Multi-View Classification

### v1

@@ Introduction
Given a feature matrix $X^{v}\in\mathbb{R}^{d_{v}\times n}$, where $d_{v}$ represents the feature dimension for view $v$ and $n$ represents the number of samples, we first compute the pairwise distance matrix $D\in\mathbb{R}^{n\times n}$, where each element $D_{ij}=\|\boldsymbol{x}_{i}^{v}-\boldsymbol{x}_{j}^{v}\|_{2}^{2}$ represents the squared distance between samples $\boldsymbol{x}_{i}^{v}$ and $\boldsymbol{x}_{j}^{v}$. Next, we sort the distances for each sample to identify its top $k$ nearest neighbors. Let $\varphi(\boldsymbol{x}_{i}^{v})$ denote the distance to the $k$-th nearest neighbor for sample $\boldsymbol{x}_{i}^{v}$. We then compute a weight matrix $W$ using the formula $T_{ij}=\varphi(\boldsymbol{x}_{i}^{v})-D_{ij}$, which balances the influence of each neighbor as follows:
$W_{ij}=\max\left(0,\frac{T_{ij}}{k\cdot\varphi(\boldsymbol{x}_{i}^{v})-\sum_{j\in N_{i}}D_{ij}}\right),$ (2)
@@
$\mathcal{L}_{acc}(\boldsymbol{\alpha}_{n}^{v})=\mathcal{L}_{ace}(\boldsymbol{\alpha}_{n}^{v})\ +\ \lambda_{t}\mathcal{L}_{KL}(\boldsymbol{\alpha}_{n}^{v}),$ (16)
where $\lambda_{t}=\min(1.0,t/\mathcal{T})\in\left[0,1\right]$ is the annealing coefficient, with $t$ representing the current training epoch and $\mathcal{T}$ the total number of annealing steps. This gradual increase in the influence of KL divergence in the loss function helps prevent premature convergence of misclassified instances towards a uniform distribution.
@@
$\vspace{-2mm}\mathcal{L}=\frac{1}{V}\sum_{v=1}^{V}\mathcal{L}_{acc}(\boldsymbol{\alpha}_{n}^{v})+\mathcal{L}_{\text{con}}$ (18)
@@
We evaluated TMDL-OA, RCML, and our TUNED model on the HandWritten dataset with varying numbers of noisy views in Figure 4. Conflict test sets were created by adding noise to single, paired, and multiple views among the six available. Results show that as more views are corrupted, accuracy generally decreases. However, our TUNED model consistently outperforms others, demonstrating lower variance and greater robustness, especially in scenarios with noise in single views, where RCML and TMDL-OA show significant instability.
@@ Conclusion
The proposed Selective Markov Random Field (S-MRF) fusion framework addresses key limitations inherent in traditional Dempster-Shafer (D-S) combination methods, particularly as seen in approaches like RCML (Xu et al. 2024).
Challenges with D-S fusion in RCML
In the RCML framework, it was established that fusing two opinions ($\omega$) could be mathematically represented as averaging two pieces of evidence, using the following formula:
@@
While earlier EDL-based methods like RCML (Xu et al. 2024) and TMDL-OA (Liu et al. 2022) attempted to address these issues by incorporating intricate loss functions and hand-crafted conflict degree measures, these approaches require careful design and hyperparameters tuning, which can be both complex and error-prone.
@@
Conflict View(s) TUNED RCML TMDL-OA

### v3

@@ Introduction
Given a feature matrix $X^{v}\in\mathbb{R}^{d_{v}\times n}$, where $d_{v}$ represents the feature dimension for view $v$ and $n$ represents the number of samples, we first compute the pairwise distance matrix $D\in\mathbb{R}^{n\times n}$, where each element $D_{ij}=\|\boldsymbol{x}_{i}^{v}-\boldsymbol{x}_{j}^{v}\|_{2}^{2}$ represents the squared distance between samples $\boldsymbol{x}_{i}^{v}$ and $\boldsymbol{x}_{j}^{v}$. Next, we sort the distances for each sample to identify its top $k$ nearest neighbors. Let $\varphi(\boldsymbol{x}_{i}^{v})$ denote the distance to the $k$-th nearest neighbor for sample $\boldsymbol{x}_{i}^{v}$. We then compute a weight matrix $\boldsymbol{\mathcal{M}}$ using the formula $T_{ij}=\varphi(\boldsymbol{x}_{i}^{v})-D_{ij}$, which balances the influence of each neighbor as follows:
$\mathcal{M}_{ij}=\max\left(0,\frac{T_{ij}}{k\cdot\varphi(\boldsymbol{x}_{i}^{v})-\sum_{j\in N_{i}}D_{ij}}\right),$ (2)
@@
$\mathcal{L}_{acc}(\boldsymbol{\alpha}_{n}^{v})=\mathcal{L}_{ace}(\boldsymbol{\alpha}_{n}^{v})\ +\ \lambda_{s}\mathcal{L}_{KL}(\boldsymbol{\alpha}_{n}^{v}),$ (16)
where $\lambda_{s}=\min(1.0,s/\mathcal{T})\in\left[0,1\right]$ is the annealing coefficient, with $t$ representing the current training epoch and $\mathcal{T}$ the total number of annealing steps. This gradual increase in the influence of KL divergence in the loss function helps prevent premature convergence of misclassified instances towards a uniform distribution.
@@
$\vspace{-2mm}\mathcal{L}=\frac{1}{V}\sum_{v=1}^{V}\mathcal{L}_{acc}(\boldsymbol{\alpha}_{n}^{v})+\lambda_{t}\mathcal{L}_{\text{con}}$ (18)
@@
We evaluated TMDL-OA, ECML, and our TUNED model on the HandWritten dataset with varying numbers of noisy views in Figure 4. Conflict test sets were created by adding noise to single, paired, and multiple views among the six available. Results show that as more views are corrupted, accuracy generally decreases. However, our TUNED model consistently outperforms others, demonstrating lower variance and greater robustness, especially in scenarios with noise in single views, where ECML and TMDL-OA show significant instability.
@@ Conclusion
The proposed Selective Markov Random Field (S-MRF) fusion framework addresses key limitations inherent in traditional Dempster-Shafer (D-S) combination methods, particularly as seen in approaches like ECML (Xu et al. 2024).
Challenges with D-S fusion in ECML
In the ECML framework, it was established that fusing two opinions ($\omega$) could be mathematically represented as averaging two pieces of evidence, using the following formula:
@@
While earlier EDL-based methods like ECML (Xu et al. 2024) and TMDL-OA (Liu et al. 2022) attempted to address these issues by incorporating intricate loss functions and hand-crafted conflict degree measures, these approaches require careful design and hyperparameters tuning, which can be both complex and error-prone.
@@
Conflict View(s) TUNED ECML TMDL-OA